Volcán Barva (04.01.2015 to 06.30.2015)

You are here

PDF versionPDF version
Year: 
2015
Quarter: 
April - June
Communications activities: 

The Site Manager attended to the official signature of the agreement of OEFO, being TEAM project is part of it. This observatory is part of the national strategy of monitoring and our ten 1ha permanent plots belong to this.

In the next weeks OEFO is planning arrange a meeting with SINAC (National System of Protected Areas – Environment authority) in order to link OEFO with the National political environment through the environmental authorities.

 

 

The site Manager attended the official presentation of the results of database of people working with camera trapping. In this workshop the National System of Protected Areas – Environment authority, pose a great opportunity to implement the National Network of Camera trapping monitoring. The site manager offered here services to write the agreement following the requirements of the Costa Rican Program of Protected areas Monitoring and Biological Corridors (PROMEC-CR).

This approach will be very useful due the incorporation of the camera trapping results and recommendations into the management and operative national plans of SINAC.

Unusual Events at the TEAM Site: 

All month of June have showed heavy rainning. The precipitation during this season have been very inusual. The base camp have been flooded, bridges collapsed and bad roads.

New Species: 

No new species at this samplig period

Protocol Activities: 

The annual census of trees, palms and lianas ≥10 cm diameter in the 1 ha plot located at 50, 600,1500, 2000, masl was completed up to date. Our annual re-census in our ten 1ha vegetation plots have been successfully completed. More than 5000 stems have been measured.

 

We have completed the terrestrial vertebrates monitoring for our sampling dry period in array 1, 2 and 3. (low middle and high elevation). 60 camera traps were deployed. Interesting and new sightings of mammals specially tapirs and deers weere recorded.  3000 pictures were processed on the deskteam. All data from Array 3 were already uploaded in the server.

 

The climate data from April and June were collected and submitted in the new data entry form version 3.0.  But we have had several problems with the station due the bad weather in the region and sometimes the energy system, datalogger and some sensors are malfunctioning.

Protocol Problems: 

No protocol problems

Schedule Problems: 

No schedule problems at this sampling period.

Logistical Problems: 

We have some logistical problems for carrying out the climate sampling due bad roads, motorcycle damage, etc due heavy rain.

Suggestions: 

I suggest some improvemnts in the webpage for climate protocol. The data status report only appears up to year 2013.

The current sensors does not appear in climate maintenance, the way to update the sensors management is confused.

I have uploaded all information required for the new sensors and the old sensors removed for calibration, but the new sensor does not appear on the climate maintenance webpage. I mean, when I am doing the climate maintenance for these Rh/T sensor, the serial number does not appear.